Suppose that three months have passed, so it is August 15, 2000. You are given the following data:

What is the value of the forward agreement? Round to three digits after the decimal. Remember that the notional is $100

Maturity Date

Z(0,T)

Nov 2000 (3 months)

0.9844

Feb 2001 (6 months)

0.9690

May 2001 (9 months)

0.9531

Aug 2001 (12 months)

0.9386
Answer #1

Answer-

Given

Date is August 15th 2000

Therefore

Z 90 = 0.9844
Z 180 = 0.9690
Z 270 = 0.9531
Z 360 = 0.9386

The quarterly fixed rate = 1 - Z 360 / ( Z 90 + Z 180 + Z 270 + Z 360)

The quarterly fixed rate  = 1 - 0.9386 / ( 0.9844 + 0.9690 + 0.9531 + 0.9386 )

The quarterly fixed rate  = 0.0614 / 3.8451

The quarterly fixed rate = 0.01597

The quarterly fixed rate  = 1.597 %

Given

The notional principal = $ 100
The quarterly payments = $ 100 x 0.01597 = $ 1.597 [ Value of Forward agreement quarterly]

The rate in annual terms = 1.597 % x (360 / 90) = 1.597 % x 4 = 6.388 % = 0.06388
The annual payments = $ 100 x 0.06388 = $ 6.388 [ Value of forward agreement annually ]